# Diffusion posterior sampling for simulation-based inference in tall data settings

### 11 Apr 2024 v1

@@ Abstract
Determining which parameters of a non-linear model could best describe a set of experimental data is a fundamental problem in science and it has gained much traction lately with the rise of complex large-scale simulators (a.k.a. black-box simulators). The likelihood of such models is typically intractable, which is why classical MCMC methods can not be used. Simulation-based inference (SBI) stands out in this context by only requiring a dataset of simulations to train deep generative models capable of approximating the posterior distribution that relates input parameters to a given observation. In this work, we consider a tall data extension in which multiple observations are available and one wishes to leverage their shared information to better infer the parameters of the model. The method we propose is built upon recent developments from the flourishing score-based diffusion literature and allows us to estimate the tall data posterior distribution simply using information from the score network trained on individual observations. We compare our method to recently proposed competing approaches on various numerical experiments and demonstrate its superiority in terms of numerical stability and computational cost.
@@ 1 Introduction
Inverting non-linear models describing natural phenomena is a fundamental problem in science and has been tackled by many fields (Gonçalves et al.,, 2020; Vasist et al.,, 2023; Dax et al.,, 2023). In this work, we consider a Bayesian approach (Tarantola,, 2005) in which the input parameters $\theta\in\mathbb{R}^{m}$ of a model $\mathcal{M}$ that best explain a set of output observations $x\in\mathbb{R}^{d}$ are described via a posterior distribution $p(\theta\mid x)$. Obtaining samples of the posterior can, however, be very challenging when the outputs of $\mathcal{M}$ are obtained through complex simulations (e.g. solutions of non-linear differential equations (Jansen and Rit,, 1995)). Indeed, in these cases the likelihood function $p(x\mid\theta)$ is often impossible to evaluate and MCMC procedures can not be used.
Simulation-based inference (SBI) (Cranmer et al.,, 2020) is a promising approach that bypasses the difficulties of likelihood evaluations via simulations from the model and leverages the recent advances from deep generative learning. The procedure relies on the choice of a prior distribution $\lambda(\theta)$ encoding knowledge about the values of $\mathcal{M}$’s parameters and a simulated dataset generated as per:
@@
Letting $N_{s}$ be a fixed simulation budget, one can build a dataset $\mathcal{D}=\{(\Theta_{i},X_{i})\}_{i=1}^{N_{s}}$ and train conditional deep generative models to generate samples from $p(\theta\mid x^{\star})$ for any new observation $x^{\star}$. Usual approaches approximate either directly the posterior distribution (NPE) (Greenberg et al.,, 2019), the likelihood function (NLE) (Papamakarios et al.,, 2019), or the ratio of likelihoods (NRE) (Hermans et al.,, 2020). In NPE, the posterior samples can be obtained directly by sampling a conditional normalizing flow (Papamakarios et al.,, 2021), whereas in NLE and NRE it is necessary to use a MCMC sampler.
In this work, we consider a natural extension of the above Bayesian framework to a tall data setting (Bardenet et al.,, 2015), in which multiple observations $x^{\star}_{1:n}=(x_{1},\dots,x_{n})$ are available and the posterior distribution
@@
is expected to provide more precise information about how to invert $\mathcal{M}$ (see Fig 1). Despite being a setup with much practical interest, there has not been many previous works providing a satisfactory extension to usual SBI procedures. In Rodrigues et al., (2021), the authors merge a fixed number of extra observations via a Deepset (Zaheer et al.,, 2017) and fall back to NPE with an augmented training dataset with samples $\mathcal{D}_{n}=\{(\Theta_{i},X_{i,1:n})\}$. Important drawbacks of such approach are the lack of flexibility on the number of extra observations at inference time and the potentially heavy cost of generating extra observations from the simulator. In Hermans et al., (2020), the authors show how to handle the tall data setting with NRE and an equivalent extension can be done for NLE as well (Geffner et al.,, 2023). Note, however, that both approaches still require a MCMC sampler to obtain posterior samples, which is often undesirable in SBI applications.
@@
Figure 1: The posterior distribution of a Gaussian simulator model with Gaussian prior concentrates around the true parameter $\theta^{\star}$ as the number $n$ of observations ${x^{\star}_{i}\sim p(x\mid\theta^{\star})}$ increases. We compare the analytic posterior to the posterior estimated with our score-based proposal (GAUSS).
@@
Recently, Sharrock et al., (2022) proposed to use score-based generative models (SBGM) (Ho et al.,, 2020; Song et al., 2021b, ) to approximate the conditional posterior distribution targeted in SBI, called Neural Posterior Score estimation (NPSE). SBGM introduces an easy-to-train objective for learning the score of a sequence of perturbed versions of the target distribution. This relies on a network $s_{\phi}(\theta,x_{i},t)$ trained using the simulated dataset. SBGM rivals state-of-the-art generative models such as generative adversarial networks (GANs) (Goodfellow et al.,, 2014) and normalizing flows in high-dimension challenging datasets without the need of adversarial training or special network architectures.
@@
However, extending NPSE to a tall data setting is challenging, as a direct application of the method would require training a score network $s_{\phi}(\theta,x_{1:n},t)$ to approximate $\nabla_{\theta}\log p_{t}(\theta\mid x_{1:n})$ on an augmented dataset $\mathcal{D}_{n}$, leading to the same drawbacks from the augmented NPE approach. Recently, Geffner et al., (2023) proposed a method named F-NPSE in which the target tall posterior is approximated via an annealed Langevin procedure, using a sequence of probability distributions for which the score can be computed using the individual scores $s_{\phi}(\theta,x_{i},t)$, thus avoiding the requirement of an augmented training dataset. The major drawback of this approach, however, is reintroducing the need of MCMC for sampling from the posterior.
@@
In this paper, we propose an algorithm that approximates the score of the tall data posteriors using only the individual scores from the individual SBGMs and without a Langevin sampler. We demonstrate the superiority of our approach as compared to F-NSPE on several numerical experiments: two Gaussian toy models for which all quantities of interest are known analytically, three examples from the SBI benchmark (Lueckmann et al., 2021a, ), and the inversion of a challenging model from neuroscience (Jansen and Rit,, 1995).
@@ 2 Background
2.1 Score based generative models (SBGM)
Score estimation.
The goal of score-based generative modelling is to learn how to sample new data from a $m$-dimensional target distribution $q_{\operatorname{data}}$ using an approximation of the score on “noisy versions” of samples from $q_{\operatorname{data}}$. In the variance preserving (VP) framework (Ho et al.,, 2020; Yang et al.,, 2023), a sequence of noisy versions $\{\theta_{t}\}_{t\in[1:T]}$ of $q_{\operatorname{data}}$ is defined for $t\in[1:T]$, by
@@ 2.1 Score based generative models (SBGM)
where $\theta_{0}\sim q_{\operatorname{data}}$, $\{\epsilon_{t}\}_{t\in[1:T]}$ are i.i.d. with distribution $\mathcal{N}(0,\mathbf{I}_{m})$, $\{\alpha_{t}\}_{t\in[1:T]}\in[0,1]^{T}$ is a decreasing sequence of positive real numbers. Let $q_{t|0}(\theta_{t}|\theta_{0})=\mathcal{N}(\theta_{t};\sqrt{\alpha_{t}}\theta_{0},(1-\alpha_{t})\mathbf{I}_{m})$ be the probability density function of the conditional distribution of $\theta_{t}$ given $\theta_{0}$ and ${{\mathsf{q}}}_{t}{}$ the density of $\theta_{t}$. Following the seminal works of Hyvärinen and Dayan, (2005); Vincent, (2011), we can learn the score of ${{\mathsf{q}}}_{t}$ via a neural network $\mathrm{s}_{\psi}$ by minimising $\mathcal{L}:\psi\mapsto\mathbb{E}_{\theta_{t}\sim{{\mathsf{q}}}_{t}}\left[\|\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})-\nabla\log{{\mathsf{q}}}_{t}(\theta_{t})\|^{2}\right]$ without knowledge of the ground-truth data score. Using Fisher’s identity, we can define the SBGM loss as
$\mathcal{L}_{\mathrm{score}}(\psi)=\sum_{t=1}^{T}\gamma_{t}^{2}\mathbb{E}_{\theta_{0}\sim q_{\operatorname{data}},\theta_{t}\sim q_{t|0}(\cdot|\theta_{0})}\Big{[}\|\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})\\ -\nabla\log q_{t|0}(\theta_{t}|\theta_{0})\|^{2}\Big{]}\,,$ (2)
@@
Backward sampling.
Once we have the score approximation $\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})$ of $\nabla\log{{\mathsf{q}}}_{t}(\theta_{t})$, our goal is to sample backwards from the distributions ${{\mathsf{q}}}_{T},\ldots,{{\mathsf{q}}}_{1}$. There are many ways of carrying out this sampling, such as using annealed Langevin dynamics (Song and Ermon,, 2019), stochastic differential equations (Song et al., 2021b, ), or ordinary differential equations (Karras et al.,, 2022). In this work, we follow the approach proposed in Song et al., 2021a , which yields the denoising diffusion implicit models (DDIM) sampler. DDIM introduces a set of inference distributions indexed by $\sigma=\{\sigma_{t}\in(0,\alpha_{t-1}^{1/2})\}_{t\in[1:T-1]}$ defined for $t\in[1:T-1]$ as:
$q^{\sigma}_{t|t+1,0}(\theta_{t}|\theta_{0},\theta_{t+1})=\mathcal{N}\left(\theta_{t-1};\bm{\mu}_{t}(\theta_{0},\theta_{t}),\sigma_{t}^{2}\mathbf{I}_{m}\right)\,,$
with $\bm{\mu}_{t}(\theta_{0},\theta_{t})=\alpha_{t-1}{}^{1/2}\theta_{0}+(\upsilon_{t-1}-\sigma^{2}_{t})^{1/2}(\theta_{t}-\alpha_{t}{}^{1/2}\theta_{0})\big{/}(1-\alpha_{t})^{1/2}$. Note that $\bm{\mu}_{t}$ is chosen so that $q^{\sigma}_{t|0}(\theta_{t}|\theta_{0})=\mathcal{N}(\theta_{T};\sqrt{\alpha_{t}}\theta_{0},(1-\alpha_{t})\mathbf{I}_{m})$  (Song et al., 2021a, , Lemma 1, Appendix B). This property allows us to write
${{\mathsf{q}}}_{t-1}(\theta_{t-1})=\int q^{\sigma}_{t-1|t,0}(\theta_{t-1}|\theta_{t},\theta_{0})\times q_{0|t}(\theta_{0}|\theta_{t}){{\mathsf{q}}}_{t}(\theta_{t})\mathrm{d}\theta_{0}\mathrm{d}\theta_{t}\,.$
Even though the equation above suggests a way of passing from ${{\mathsf{q}}}_{t}$ to ${{\mathsf{q}}}_{t-1}$ it involves an intractable kernel $\int q^{\sigma}_{t-1|t,0}(\theta_{t-1}|\theta_{t},\theta_{0})q_{0|t}(\theta_{0}|\theta_{t})\mathrm{d}\theta_{0}$. The marginal distribution can be approximated by
@@
where $\mu_{t}(\theta_{t})$ is the conditional mean of $\theta_{0}$ given $\theta_{t}$ under the inference distribution; i.e. $\mu_{t}(\theta_{t})\vcentcolon=\mathbb{E}_{\theta_{0}\sim q^{\sigma}_{0|t}(\cdot|\theta_{t})}\left[\theta_{0}\right]$. As $\mu_{t}(\theta_{t})$ is not available explicitly, in order to effectively sample from (3), we can estimate $\mu_{t}(\theta_{t})$ using the approximate score $\mathrm{s}_{\psi}(\theta_{t},t)$. By noting that
$\alpha_{t}^{1/2}\mu_{t}(\theta_{t})-\theta_{t}=(1-\alpha_{t})\mathbb{E}_{\theta_{0}\sim q^{\sigma}_{0|t}(\cdot|\theta_{t})}\left[\nabla\log q_{t|0}(\theta_{t}|\theta_{0})\right]\,,$
we can use (2) to define the following approximation for all $t\in[1:T]$
$\mu_{\psi,t}(\theta_{t})\vcentcolon=\alpha_{t}^{1/2}\Big{(}\theta_{t}+(1-\alpha_{t})\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})\Big{)}\,.$ (4)
We can finally define the backward Markov chain
@@
Note that we use the procedure described above to learn the score of the conditional distribution of $\theta$ given $x$. To do so, we learn an amortized score $\mathrm{s}_{\psi}(\theta,x,\alpha_{t})$ as in Batzolis et al., (2021).
2.2 Factorized neural posterior score estimation
The F-NPSE method proposed in Geffner et al., (2023) defines a sequence of distributions $\{\varrho_{t}(\cdot\mid x)\}_{t\in[0:T]}$ as per
$\nabla_{\theta}\log\varrho_{t}(\theta\mid x^{\star}_{1:n})=(1-n)(1-t)\nabla_{\theta}\log\lambda(\theta)\\ +\sum_{j=1}^{n}\mathrm{s}_{\psi}(\theta,x_{i},\alpha_{t})\,,$ (6)
where the score of $\varrho_{0}$ is the score of the tall posterior defined in (1) and $\nabla_{\theta}\log\lambda(\theta)$ can be computed analytically for common prior choices; see Appendix D. F-NPSE uses Langevin dynamics to sample $\varrho_{t}$ for each $t=T,\ldots,1$ and has shown superior performance as compared to all competing methods for tall data settings (i.e. augmented versions of NPE, NLE, and NRE) in terms of posterior reconstruction, achieving the best trade-off between sample efficiency and accumulation of errors when the number of observations grows. However, this performance is at the cost of an increased number of neural net evaluations. The main limitation of F-NPSE is the use of annealed Langevin dynamics which is very sensitive to the choices of step-size at each sampling iteration, as well as the total number of steps. We justify these claims with empirical results in Section 4.1 and Appendix J.

### 07 Jun 2024 v2

@@ Abstract
Determining which parameters of a non-linear model best describe a set of experimental data is a fundamental problem in science and it has gained much traction lately with the rise of complex large-scale simulators. The likelihood of such models is typically intractable, which is why classical MCMC methods can not be used. Simulation-based inference (SBI) stands out in this context by only requiring a dataset of simulations to train deep generative models capable of approximating the posterior distribution that relates input parameters to a given observation. In this work, we consider a tall data extension in which multiple observations are available to better infer the parameters of the model. The proposed method is built upon recent developments from the flourishing score-based diffusion literature and allows to estimate the tall data posterior distribution, while simply using information from a score network trained for a single context observation. We compare our method to recently proposed competing approaches on various numerical experiments and demonstrate its superiority in terms of numerical stability and computational cost.
@@ 1 Introduction
Inverting non-linear models describing natural phenomena is a fundamental problem in science and has been tackled by many fields (Gonçalves et al.,, 2020; Dax et al.,, 2023). In this work, we consider a Bayesian approach in which the input parameters $\theta\in\mathbb{R}^{m}$ of a model $\mathcal{M}$ that best explain a set of output observations $x\in\mathbb{R}^{d}$ are described via a posterior distribution $p(\theta\mid x)$. Obtaining samples of the posterior can, however, be very challenging when the outputs of $\mathcal{M}$ are obtained through complex simulations (e.g. solutions of non-linear differential equations). Indeed, in these cases the likelihood $p(x\mid\theta)$ is often impossible to evaluate and MCMC procedures cannot be used.
Simulation-based inference (SBI) (Cranmer et al.,, 2020) is a promising approach that bypasses the difficulties of likelihood evaluations via simulations from the model and leverages the recent advances from deep generative learning. The procedure relies on the choice of a prior distribution $\lambda(\theta)$ encoding knowledge about the values of $\mathcal{M}$’s parameters and a simulated dataset generated as:
@@
Let $N_{s}$ be a fixed simulation budget, we can sample a dataset $\mathcal{D}=\{(\Theta_{i},X_{i})\}_{i=1}^{N_{s}}$ and train conditional deep generative models to generate samples from $p(\theta\mid x^{\star})$ for any new observation $x^{\star}$. Usual approaches target either the posterior directly (NPE) (Greenberg et al.,, 2019), the likelihood function (NLE) (Papamakarios et al.,, 2019), or the likelihood-to-evidence ratio (NRE) (Hermans et al.,, 2020). In NPE, the posterior samples can be obtained directly by sampling a conditional normalizing flow (Papamakarios et al.,, 2021), whereas in NLE and NRE additional MCMC sampling is required.
In this work, we consider a natural extension of the above Bayesian framework to the tall data setting (Bardenet et al.,, 2015), in which multiple observations $x^{\star}_{1:n}=(x_{1},\dots,x_{n})$ are available and
@@
the tall data posterior, is expected to provide more precise information about how to invert $\mathcal{M}$ (see Fig 1). While this framework is crucial for practical applications, there is no satisfactory solution to process such a tall data setting in SBI procedures. In Rodrigues et al., (2021), the authors merge a fixed number of extra observations via a Deepset (Zaheer et al.,, 2017) and fall back to NPE trained on an augmented dataset with samples $\mathcal{D}_{n}=\{(\Theta_{i},X_{i,1:n})\}$. Important drawbacks of this approach are the potentially heavy cost of generating extra observations via the simulator ($N_{s}^{\mathrm{aug}}=N_{s}\times n$) and the lack of flexibility on the number of extra observations at inference time. In Hermans et al., (2020), the authors show how the amortization of NRE allows to handle the tall data setting without requiring new simulations or retraining, by factorizing over the set of multiple observations. An equivalent extension can be done for NLE (Geffner et al.,, 2023). Note, however, that both approaches still require a MCMC sampler to obtain posterior samples, which is often undesirable in SBI applications.
@@
Figure 1: The posterior distribution of a model with a Gaussian simulator and Gaussian prior concentrates around the true parameter $\theta^{\star}$ as the number $n$ of observations ${x_{i}\sim p(x\mid\theta^{\star})}$ increases. In this simple setting we can calculate the analytic posterior distribution and compare it to the posterior estimated with our score-based proposal (GAUSS).
@@
Recently, Sharrock et al., (2022) proposed a new method that relies on score-based generative modelling (SBGM) (Ho et al.,, 2020; Song et al., 2021b, ) to approximate the posterior distribution targeted in SBI, called Neural Posterior Score Estimation (NPSE). It introduces an easy-to-train objective to learn the score $\nabla_{\theta}\log p_{t}(\theta\mid x)$ of a sequence of noisy versions of the target distribution, using the simulated dataset $\mathcal{D}$. Sampling from the posterior then consists in progressively denoising the perturbed data by solving the backward diffusion process using the learned score. SBGM rivals state-of-the-art generative models such as generative adversarial networks (GANs) (Goodfellow et al.,, 2014) and normalizing flows (Papamakarios et al.,, 2021) in challenging high-dimensional datasets, without the need of adversarial training or special (e.g. invertable) network architectures. Indeed, NPSE trained with transformer architectures is currently the most flexible SBI algorithm and achieves state-of-the-art performance on benchmark tasks, as recently shown by Gloeckler et al., (2024).
@@
However, extending NPSE to the tall data setting is challenging, as a direct application of the method would require training a score network $s_{\phi}(\theta,x_{1:n},t)$ to approximate $\nabla_{\theta}\log p_{t}(\theta\mid x_{1:n})$ on an augmented dataset $\mathcal{D}_{n}$, which leads to the same drawbacks as for the augmented NPE approach. Geffner et al., (2023) recently proposed F-NPSE that, like Hermans et al., (2020) for NRE, exploits the amortization of NPSE to approximate the tall data posterior without the need of an augmented dataset or retraining the score network. To do so, F-NPSE introduces a sequence of distributions for which the score can be computed by composing individual scores $s_{\phi}(\theta,x_{j},t)$ obtained for each observation $x_{j}$, following the factorization from Equation (1). The major drawback of this approach, however, is that the diffusion process followed by this sequence is unknown, which reintroduces the need of MCMC methods to sample from the posterior via an annealed Langevin proceedure.
@@
In this paper, we propose a new sampling algorithm that approximates the tall data posterior using only the individual scores obtained from an amortized NPSE, but without the need of Langevin dynamics, by directly solving the backward diffusion process for the tall posterior using the factorization from Equation (1). This sampler relies on a second order approximation of the backward diffusion kernels associated with the individual posterior scores. We demonstrate the superiority of our approach as compared to F-NPSE on several numerical experiments: two Gaussian toy models for which all quantities of interest are known analytically, three examples from the SBI benchmark, and the inversion of a challenging model from computational neuroscience.
@@ 2 Background
Score based generative models (SBGM).
The goal of SBGM is to estimate an unknown target data distribution $q_{\operatorname{data}}$ from samples $\theta_{1},\dots,\theta_{M}\in\mathbb{R}^{d}\sim_{\mathrm{iid}}q_{\operatorname{data}}$ with the help of a forward diffusion process that adds noise to the training data. The main idea is to approximately sample the target distribution by solving the associated backward diffusion process. This procedure requires to approximate the score functions of the diffused data distributions for different levels of added noise.
In this work, we focus on the variance preserving (VP) framework (Ho et al.,, 2020; Yang et al.,, 2023) in which a sequence of noisy versions $\{\theta_{t}\}_{t\in[1:T]}$ of $q_{\operatorname{data}}$ is defined for $t\in[1:T]$, by
@@
where $\theta_{0}\sim q_{\operatorname{data}}$, $\{\epsilon_{t}\}_{t\in[1:T]}$ are i.i.d. with distribution $\mathcal{N}(0,\mathbf{I}_{m})$ and $\{\alpha_{t}\}_{t\in[1:T]}\in[0,1]^{T}$ is a decreasing sequence of positive real numbers. Let $q_{t|0}(\theta_{t}|\theta_{0})=\mathcal{N}(\theta_{t};\sqrt{\alpha_{t}}\theta_{0},(1-\alpha_{t})\mathbf{I}_{m})$ be the probability density function of the conditional distribution of $\theta_{t}$ given $\theta_{0}$ and ${{\mathsf{q}}}_{t}{}$ the density of $\theta_{t}$. Following Hyvärinen and Dayan, (2005); Vincent, (2011), we can learn the score of ${{\mathsf{q}}}_{t}$ via a neural network $\mathrm{s}_{\psi}$ by minimising $\mathcal{L}:\psi\mapsto\mathbb{E}_{\theta_{t}\sim{{\mathsf{q}}}_{t}}\left[\|\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})-\nabla\log{{\mathsf{q}}}_{t}(\theta_{t})\|^{2}\right]$ without knowledge of the ground-truth data score. Using Fisher’s identity, we can define the denoising score-matching loss
$\mathcal{L}_{\mathrm{score}}(\psi)=\sum_{t=1}^{T}\gamma_{t}^{2}\mathbb{E}_{\theta_{0}\sim q_{\operatorname{data}},\theta_{t}\sim q_{t|0}(\cdot|\theta_{0})}\Big{[}\|\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})-\nabla\log q_{t|0}(\theta_{t}|\theta_{0})\|^{2}\Big{]}\,,$ (2)
@@
Once we have the score approximation $\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})$ of $\nabla\log{{\mathsf{q}}}_{t}(\theta_{t})$, the goal is to draw backwards starting from $\theta_{T}\sim{{\mathsf{q}}}_{T}$ to obtain samples approximately distributed according to $q_{\operatorname{data}}$ at time 0. Different ways to perform this backward sampling include the use of annealed Langevin dynamics (Song and Ermon,, 2019), stochastic differential equations (Song et al., 2021b, ), or ordinary differential equations (Karras et al.,, 2022). In this work, we follow the approach proposed in Song et al., 2021a , which yields the denoising diffusion implicit models (DDIM) sampler. DDIM introduces a set of inference distributions indexed by $\sigma=\{\sigma_{t}\in(0,\alpha_{t-1}^{1/2})\}_{t\in[1:T-1]}$ defined for $t\in[1:T-1]$ as $q^{\sigma}_{t|t+1,0}(\theta_{t}|\theta_{0},\theta_{t+1})=\mathcal{N}\left(\theta_{t-1};\bm{\mu}_{t}(\theta_{0},\theta_{t}),\sigma_{t}^{2}\mathbf{I}_{m}\right)$, with $\bm{\mu}_{t}(\theta_{0},\theta_{t})=\alpha_{t-1}{}^{1/2}\theta_{0}+(\upsilon_{t-1}-\sigma^{2}_{t})^{1/2}(\theta_{t}-\alpha_{t}{}^{1/2}\theta_{0})\big{/}(1-\alpha_{t})^{1/2}$. Note that $\bm{\mu}_{t}$ is chosen so that $q^{\sigma}_{t|0}(\theta_{t}|\theta_{0})=\mathcal{N}(\theta_{t};\sqrt{\alpha_{t}}\theta_{0},(1-\alpha_{t})\mathbf{I}_{m})$  (Song et al., 2021a, , Lemma 1, Appendix B). This property allows us to write ${{\mathsf{q}}}_{t-1}(\theta_{t-1})=\int q^{\sigma}_{t-1|t,0}(\theta_{t-1}|\theta_{t},\theta_{0})q_{0|t}(\theta_{0}|\theta_{t}){{\mathsf{q}}}_{t}(\theta_{t})\mathrm{d}\theta_{0}\mathrm{d}\theta_{t}$. Even though the equation above suggests a way of passing from ${{\mathsf{q}}}_{t}$ to ${{\mathsf{q}}}_{t-1}$, it involves an intractable kernel $\int q^{\sigma}_{t-1|t,0}(\theta_{t-1}|\theta_{t},\theta_{0})q_{0|t}(\theta_{0}|\theta_{t})\mathrm{d}\theta_{0}$. The marginal distribution can be approximated by
@@
where $\mu_{t}(\theta_{t}):=\mathbb{E}_{\theta_{0}\sim q^{\sigma}_{0|t}(\cdot|\theta_{t})}\left[\theta_{0}\right]$, and verifies
$\alpha_{t}^{1/2}\mu_{t}(\theta_{t})-\theta_{t}=(1-\alpha_{t})\mathbb{E}_{\theta_{0}\sim q^{\sigma}_{0|t}(\cdot|\theta_{t})}\left[\nabla\log q_{t|0}(\theta_{t}|\theta_{0})\right]\,.$
Using (2), we can now define the following approximation of $\mu_{t}(\theta_{t})$ for all $t\in[1:T]$
$\mu_{\psi,t}(\theta_{t}):=\alpha_{t}^{1/2}\Big{(}\theta_{t}+(1-\alpha_{t})\mathrm{s}_{\psi}(\theta_{t},\alpha_{t})\Big{)}\,.$ (4)
We finally obtain the following backward Markov chain for DDIM:
@@
Factorized neural posterior score estimation.
The F-NPSE method proposed in Geffner et al., (2023) defines a sequence of distributions $\{\varrho_{t}(\cdot\mid x)\}_{t\in[0:T]}$ such that
$\nabla_{\theta}\log\varrho_{t}(\theta\mid x^{\star}_{1:n})=(1-n)(1-t)\nabla_{\theta}\log\lambda(\theta)+\sum_{j=1}^{n}\mathrm{s}_{\psi}(\theta,x_{i},\alpha_{t})\,,$ (6)
where the score of $\varrho_{0}$ is the score of the tall posterior defined in (1). F-NPSE uses Langevin dynamics to sample $\varrho_{t}$ for each $t=T,\ldots,0$ and has shown superior performance as compared to all competing methods for tall data settings (i.e. augmented or factorized versions of NPE, NLE, and NRE) in terms of posterior reconstruction, achieving the best trade-off between sample efficiency and accumulation of errors when the number of observations grows. The main limitation of F-NPSE is the use of Langevin dynamics, which can be very sensitive to the choices of step-size at each sampling iteration, as well as the total number of steps. We justify these claims with empirical results in Section 4.1 and Appendix K.